HUBBLE SPACE TELESCOPE DAILY REPORT #2247 PERIOD COVERED: 0000Z (UTC) 11/09/98 - 0000Z (UTC) 11/10/98 Daily Status Report as of 314/0000Z 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed Two Sets of NIC/1/2/3 7850 (Cosmological Parameters Omega and Lambda from High-Z Type Ia Supernovae) The Near Infrared Camera (Cameras One, Two and Three) was used to make observations of high Z type Ia Super Novae 5H1.3ZA and SN97201 in order to measure the Omega and Lambda parameters. Type Ia supernovae can be used as calibrated standard candles to measure large distances. As described in HSTAR 6766 and 2.1, the acquisition for the second iteration of this proposal defaulted to fine lock backup on one FGS only. Otherwise, the observations were completed as planned, and no further anomalies were reported. 1.2 Completed Three Sets of STIS/CCD 7948 (CCD Dark and Bias Monitor -- Cont. to Dec. 98) The Space Telescope Imaging Spectrograph (CCD) was used to measure CCD dark current and bias rates for the two supported amplifier settings {GAIN=1 and GAIN=4} in order to gather data for making superdarks and superbiases. The data will be used to track the growth of hot pixels on the CCD in general and within the various target acquisition apertures specifically. The observations were executed as scheduled, and no anomalies were noted. 1.3 Completed Five Sets of WF/PC-2 7620 (WF/PC-2 Cycle 7 Standard Darks) The WF/PC-2 was used to make a series of Cycle 7 standard dark calibration images. This dark calibration program obtains dark frames every week in order to provide data for the ongoing calibration of the CCD dark current rate, and to monitor and characterize the evolution of hot pixels. As described in yesterday's report and HSTAR 6765, re-acquisitions for the final four iterations of this proposal failed, and the take data flag remained down. Otherwise, the observations were executed as scheduled, and no further anomalies were noted. 1.4 Completed Five Sets of STIS/MA1 7950 (Cycle 7 MAMA Dark Measurements) The Space Telescope Imaging Spectrograph (MA1) was used to make a dark calibration image. This is to establish if the detectors are performing nominally, that no internal movement of the MCP has taken place, and establish dark level in orbit for a given operating temperature. These darks will be used in later image reduction and also trended over instrument life time as indicators of detector degradation or damage. As described in yesterday's report and HSTAR 6765, re-acquisitions for the first four iterations of this proposal failed, and the take data flag remained down. Otherwise, the observations were completed as planned, and no further problems were reported. 1.5 Completed WF/PC-2 7713 (WF/P-2 Cycle 7 Supplemental Darks pt3/3) The WF/PC-2 was to be used to make supplemental dark calibration images. This dark calibration program obtains three dark frames every day to provide data for monitoring and characterizing the evolution of hot pixels. As described in yesterday's report and HSTAR 6765, re-acquisitions for this proposal failed, and the take data flag remained down. 1.6 Completed Two Sets of NIC/1/2/3 7963 (Darks: Warming Up) The Near Infrared Camera (Cameras One, Two and Three) was used to monitor the detector dark current during the interval between the end of science observations and the exhaustion of cryogen and subsequent warming of the dewar to > 100K. As described in yesterday's report and HSTAR 6765, re-acquisitions for the first iteration of this proposal failed, and the take data flag remained down. Otherwise, the observations were completed as planned, and no further anomalies were noted. 1.7 Completed Two Sets of NIC/2 7226 (Search for Massive Jupiters) The Near Infrared Camera (Camera Two) was used to make a corographic search for massive planets around nearby, young main sequence stars. Because of the extreme youth of these objects, any low-mass brown dwarf and planetary companions will still be in a higher luminosity phase and thus easily detectable. A dark calibration image was also taken. As described in yesterday's report and HSTAR 6765, re-acquisitions for the first iteration of this proposal failed, and the take data flag remained down. Otherwise, the observations were executed as scheduled, and no further anomalies were noted. 1.8 Completed NIC/2 7227 (A Search for Low Mass/Sub-Luminous Companions to M-Stars) The Near Infrared Camera (Camera Two) was to be used to make a search for low mass/sub-luminous companions to low mass M class stars. Knowledge of stellar and sub-stellar masses and luminosities at and below the ~0.08 M{sun} H-burning limit is of fundamental importance. As described in yesterday's report and HSTAR 6765, re-acquisitions for this proposal failed, and the take data flag remained down. 1.9 Completed NIC/2 7329 (The Nature of the Damped LyAlpha Absorbers -- A New Study of Young Galaxies) The Near Infrared Camera (Camera Two) was used to make observations of the young galaxy Q0454+0356 in order to study the nature of its damped Lyman-Alpha Absorbers (DLAs). The DLA line systems, occasionally found in quasar spectra, provide one of only a few practical methods of detecting and studying significant numbers of galaxies at high redshifts. The observations were executed as scheduled, and no anomalies were noted. 1.10 Completed WF/PC-2 7590 (Cosmological Parameters Omega and Lambda From High-Redshift Type IA Supernovae) The WF/PC-2 was used to make observations of type 1A supernova 23H1.3ZA in order to study its redshift. The HST was used to make direct measurements of the cosmological parameters Omega, Lambda, and curvature using Type 1A supernovae as a calibrated "standard candle". The observations were executed as scheduled, and no anomalies were noted. 1.11 Completed Four Sets of STIS/CCD 8084 (POMS Test Proposal: STIS Non-Scripted Parallel Proposal Continuation) The Space Telescope Imaging Spectrograph (CCD) was used to make non-scripted parallel high galactic latitude observations as part of POMS test proposal. The observations were executed as scheduled, and no anomalies were noted. 1.12 Completed Two Sets of NIC/1/2/3 7961 (Flats: Warming Up) The Near Infrared Camera (Cameras One, Two and Three) was used to monitor the flat fields during the interval between the end of science observations and the exhaustion of cryogen and subsequent warming of the dewar to > 100K. These flats will provide a monitor for particulate comtamination and detector lateral position {from the coronagraphic spot and FDA vignetting}. The proposal completed with no reported problems. 1.13 Completed Three Sets of NIC/1/2/3 7963 (Darks: Warming Up) The Near Infrared Camera (Cameras One, Two and Three) was used to monitor the detector dark current during the interval between the end of science observations and the exhaustion of cryogen and subsequent warming of the dewar to > 100K. The observations were completed as planned, and no anomalies were noted. 1.14 Completed WF/PC-2 7619 (WF/PC-2 Cycle 7 Decontamination) The WF/PC-2 was used to conduct a Cycle 7 decontamination procedure. Ultraviolet-blocking contaminants were removed, and hot pixels cured, by warming the CCDs to +20C for 6 hours. A variety of internal images were taken before and after the decontamination in order to verify its success and to monitor the instrument health. The procedures were conducted as scheduled, and no problems were reported. 1.15 Completed WF/PC-2 7909 (POMS Test Proposal: WF II Parallel Archive Proposal) The WF/PC-2 was used to make parallel observations as part of the POMS test proposal. This test proposal is designed to simulate scientific plans. The observations were completed as planned, and no problems were reported. 1.16 Completed NIC/2 7894 (Completing A Near-Infrared Search for Very Low Mass Companions to Stars within 10 pc) The Near Infrared Camera (Camera Two) was used to make observations searching for very low mass (VLM) companions to the star CD-68D47. This search will be the largest, most sensitive, volume-limited search for VLM companions ever undertaken. The observations were completed as planned, and no anomalies were reported. 1.17 Completed NIC/1/2/3 7857 (Investigating the Missing Link Between Disks Around Pre-Main Sequence and Main Sequence Stars) The Near Infrared Camera (Cameras One, Two and Three) was used to make observations of star HD100546-N. Various observations and models of the circumstellar disk around the Main Sequence IR excess star Beta Pictors tend to show that the optically thin dust disk is the result of destruction processes among planetesimals. The observations were completed as planned, and no problems were reported. 1.18 Completed NIC/1/2/3 7893 (The Two Types of BL Lac Objects: Extrema of Jet Physics) The Near Infrared Camera (Cameras One, Two and Three) was used to make observations of the BL Lac object MS2143+070. These observations will build on the WF/PC-2 survey of a well-defined sample of 97 BL Lac objects, for which host galaxies were detected in 90 simulations predict ~100 provide essential constraints on galaxy properties -- luminosity, size, morphology, and color -- including an unprecedented ensemble of host galaxy color profiles at subarcsecond spatial resolution. The observations were executed as scheduled, and no anomalies were noted. 1.19 Completed Three Sets of NIC/2 7233 (Dust Disks Around Main Sequence Stars) The Near Infrared Camera (Camera Two) was used to make observations of the dust disk around main sequence stars. The observations were made with the coronagraph to minimize the effects of glare from the bright central star. A dark calibration image was also taken. The observations were completed as scheduled, and no anomalies were reported. 2.0 FLIGHT OPERATIONS SUMMARY 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 16 Successful : 15 Scheduled Reacquisitions : 5 Successful : 4 HSTAR 6766 describes the acquisition at 3133/132607Z which defaulted to fine lock backup on FGS-1 only. The proposal mentioned in 1.1 may have been affected. Following two FHST full maneuver updates that failed at 314/0749Z and at 314/0752Z, the following acquisiton at 314/0800Z also failed, as did the re-acquisition at 314/0916Z. HSTAR 6767 was written. Proposals described in tomorrow's report were affected: two iterations of WF/PC-2 7623 as well as 7713, and NICMOS 7233. 2.2 FHST Updates: Scheduled : 30 Successful : 28 See explanation above. 2.3 Operations Notes: The SSR EDAC error counter was cleared three times using ROP SR-1. HST operations transferred ground operations from PRS1 to PRS2 at 314/2200Z to allow for scheduled maintenance of PRS1. No problems were encountered. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Continuation of normal science observations and calibrations. /DMH